# Prompting Techniques for Reducing Social Bias in LLMs through System 1 and System 2 Cognitive Processes

### v3

@@ Abstract
Dual process theory posits that human cognition arises via two systems. System 1, which is a quick, emotional, and intuitive process, which is subject to cognitive biases, and System 2, is a slow, onerous, and deliberate process. NLP researchers often compare zero-shot prompting in LLMs to System 1 reasoning and chain-of-thought (CoT) prompting to System 2. In line with this interpretation, prior research has found that using CoT prompting in LLMs leads to reduced gender bias. We investigate the relationship between bias, CoT prompting, a debiasing prompt, and dual process theory in LLMs directly. We compare zero-shot CoT, debiasing, and a variety of dual process theory-based prompting strategies on two bias datasets spanning nine different social bias categories. We incorporate human and machine personas to determine whether the effects of dual process theory in LLMs exist independent of explicit persona models or are based on modeling human cognition. We find that a human persona, debiasing, System 2, and CoT prompting all tend to reduce social biases in LLMs, though the best combination of features depends on the exact model and bias category—resulting in up to a 19 percent drop in stereotypical judgments by an LLM.111 Datasets and code are available at https://github.com/kamruzzaman15/Reduce-Social-Bias-in-LLMs
@@ 1 Introduction
Figure 1: Example of Standard Prompting and Human Persona with System 2 Prompting for Llama 2 model in the race bias category
In recent years, large language models (LLMs) like GPT-4 (Achiam et al., 2023), ChatGPT (Brown et al., 2020), Llama 2 (Touvron et al., 2023) have revolutionized many aspects of technology and society. These models display remarkable linguistic capabilities, crafting responses that not only mimic human language but also exhibit a depth of understanding previously unattainable in automation (Karanjai and Shi, 2024). A notable advancement in enhancing the reasoning capabilities of LLMs has been the introduction of CoT prompting (Wei et al., 2022). By simulating step-by-step reasoning, CoT prompting helps LLMs achieve higher levels of clarity and accuracy in complex tasks, significantly reducing errors inherent in simpler prompt designs.
Despite these advancements, LLMs continue to struggle with embedded social biases, which raises questions regarding the ethical use of LLMs in real-life applications. These biases are difficult to identify and even more challenging to eliminate due to the complex and opaque inner workings of LLMs, the flexible and nuanced nature of human language, and the culturally dependent social rules that accompany language use. This task of mitigating social biases in LLMs is paramount to ensuring fairness and inclusivity in AI-driven communication and decisions. Applying dual process theory, a well-established psychological framework, to recent AI advancements illuminates possible pathways to enhancing the reliability and ethical footprint of LLMs by identifying where LLM generations align with and diverge from human cognitive processes.
In this paper, we use dual process theory-based prompting strategies, comparing their efficacy across multiple categories of social bias from two bias datasets. Our approach incorporates human- and machine-like personas to examine whether the effects of these cognitive theories in LLMs are dependent on explicit co-modeling of human cognition or always implicitly modeled. We follow-up on this analysis by examining interactions with debiasing prompts designed specifically for social bias reduction. Finally, we compare these results to the use of CoT prompting to test whether this prompting technique aligns with System 2 reasoning in LLMs as some have claimed in the past.
Figure 1 shows an example of how the human persona with System 2 prompting reduces stereotypical engagement over standard prompting. When we use standard zero-shot prompting in Figure 1 (a), we see that Llama 2 responds with a stereotypical answer. When we use human persona with System 2 prompting in Figure 1 (b), it instead responds with an anti-stereotypical answer.
@@
  • We find that incorporating a human persona is critical for controlling for biases in LLMs. While System 2 and explicit debiasing slightly reduce stereotypical responses on their own, combining them with a human persona lead to substantial improvements and the largest reductions in bias when averaged across models and bias categories.
  • In line with Nighojkar’s (2024) recent results in the reasoning domain, we find that CoT prompting does not behave similarly to prompts that directly model System 2 for social biases. In fact, the rate of stereotypical responses is closest between CoT and System 1 prompts in all Persona variants (none, Human, and Machine). This contradicts often stated assumptions by researchers in the past (Hagendorff et al., 2023).
@@ 2 Related Work
Recent studies have explored how reasoning in LLMs can exhibit biases similar to human cognitive processes. Hagendorff et al. (2023) look into human-like reasoning biases in LMs. They show that as these models became bigger and more complex, they began making intuitive mistakes, like those found in human System 1 thinking. This trend shifted with the introduction of ChatGPT models, which effectively avoid these reasoning traps by employing chain-of-thought processes reminiscent of human System 2 thinking, even when such explicit reasoning is inhibited. In a significant enhancement to dual-process interaction within LLMs, Lin et al. (2024) introduce SWIFTSAGE a new dual-module framework for better action planning in complex interactive reasoning tasks. This framework combines behavior cloning and prompting large language models. It includes the SWIFT module for quick, intuitive responses, and the SAGE module for careful, detailed planning. Tested on 30 tasks in the ScienceWorld benchmark, SWIFTSAGE greatly outperforms current methods like SayCan, ReAct, and Reflexion. It shows its ability to efficiently solve complicated interactive challenges with less computational needs.
Coming to the debiasing studies, Furniturewala et al. (2024) investigate the use of structured prompting techniques for debiasing language models. The study explores three categories of prompts: Prefix Prompting, Self-Refinement, and Implication Prompting. Chisca et al. (2024) focus on mitigating biases in LLMs through prompt-tuning. This approach introduces small sets of trainable token embeddings that are concatenated to input sequences, aiming to reduce biases without major modifications to the model’s architecture.
@@
Dual Process Theory is a psychological account of how human thinking and decision-making arise from two distinct modes. It distinguishes between fast, automatic (System 1), and slow, effortful (System 2) modes of thinking. System 1 enables quick comprehension through associations and pre-existing knowledge. In contrast, System 2 engages when we encounter complex or novel situations that require careful thought, evaluating logical relations, and conducting explicit reasoning to arrive at conclusions. These systems guide our reasoning, decision-making, and learning processes in various cognitive tasks (Frankish, 2010). The theory illuminates the intricate relation between intuitive, heuristic thinking and analytical, rule-based cognition (Evans and Stanovich, 2013; Ferreira and Huettig, 2023). Our understanding of our own thinking and knowing our mind’s state is connected to this two-part idea. System 1 lets us quickly guess what another person is thinking in analogy to our own while System 2 helps us to think more about their state more systematically with less self-attribution to make a metacognitive judgment (Carruthers, 2009). While the Dual Process Theory first suggested that reasoning biases come from relying too heavily on System 1 and that triggering System 2 more frequently can avoid such pitfalls in thinking, newer studies show that logic and probability can be understood intuitively as well (Ferreira and Huettig, 2023). Interestingly, biases are not only caused by System 2 not getting involved. They can also come from a fight between heuristic and logical intuitions that happen at the same time. This shows that logical thinking does not just belong to System 2 (De Neys and Pennycook, 2019). These more recent developments in the theory reveal a more nuanced picture of the ideal cognitive system selection for any given task. Bellini-Leite (2023) discusses how methods such as CoT and tree-of-thought prompting in LLMs are suggestive of System 2 human reasoning, potentially mitigating frequent errors and enhancing reliability in these models. Nighojkar (2024) go on to test this correspondence by comparing against results from human experiments. He finds that CoT prompting does not simply mimic System 2. Rather, it leads to better agreement with human responses in both System 1 and System 2 triggering instructions.
@@
Recent research on LLMs has found that assigning personas to LLMs can notably impact their reasoning and responses. Beck et al. (2024) highlights that sociodemographic prompting can significantly influence model predictions and improve zero-shot learning performance in subjective tasks. But the effectiveness of this approach varies across different models, dataset types, and sociodemographic factors. de Araujo and Roth (2024) explore how different personas, such as roles or identities (e.g., gender, occupation, or cultural background) influence the behavior and outputs of LLMs across various tasks. The research involves giving 162 unique personas to various LLMs and assessing how well they perform on different tasks and it turns out that personas not only affect task performance, bias, and model refusals but also show varied responses depending on the assigned persona. In addition to personas, explicitly debiasing instructions in prompts have been found to effectively reduce gender bias in LLMs (Kaneko et al., 2024).
@@ 3 Dataset
We use two bias detection datasets to measure the behavior of the LLMs, StereoSet (Nadeem et al., 2020) and Kamruzzaman et al.’s (Kamruzzaman et al., 2024) bias detection dataset. StereoSet covers gender, race, religion, and profession biases, specifically focusing on stereotypes related to these categories. Kamruzzaman et al.’s bias detection dataset covers age, beauty, institution, beauty in the profession, and nationality bias and how LLMs associate specific attributes in these categories to unrelated general positive and negative descriptors. We will call this dataset the GenAssocBias dataset for the remainder of this paper.
@@ 4 Experimental Setup
We evaluate five major language models in this paper: 1) GPT-4 (Achiam et al., 2023), using the GPT-4 checkpoint on the OpenAI API; [^2] Due to the cost of running GPT-4, we use a sample of 2,100 entries from the GenAssocBias dataset to run GPT-4. We ensured that this sample contains an equal number of examples from each bias category. 2) GPT-3.5 (Brown et al., 2020), using the GPT-3.5-TURBO-INSTRUCT checkpoint on the OpenAI API; 3) Llama-2-13B (Touvron et al., 2023), using the TheBloke/Llama-2-13B-chat-GGUF [^3] We use 4-bit quantized Llama 2 and Mistral models due to limited local computing resources. checkpoint on Huggingface; 4) Mistral-7B (Jiang et al., 2023), using the TheBloke/Mistral-7B-Instruct-v0.1-GGUF checkpoint on Huggingface; 5) Gemini (Team et al., 2023), using gemini-1.0-pro-latest on the Google Gemini API.
We use 12 different types of prompting techniques in the main experiments including the combinations of CoT, System 1, System 2, and Persona. This is followed up with 6 additional prompting techniques with explicit debiasing.
@@
HP + CoT Adopt the identity of a person who answers questions by thinking them “step-by-step". Fill in the BLANK while staying in strict accordance with the nature of this identity. Given the context below, …
HP + System 2 + CoT Adopt the identity of a person who answers questions slowly and thoughtfully. Their answers are effortful and reliable. They answer questions by thinking about them “step-by-step". Fill in the BLANK while staying in strict accordance with the nature of this identity. Given the context below, …
@@
MP + CoT Adopt the identity of a machine that answers questions by analyzing them “step-by-step". Fill in the BLANK while staying in strict accordance with the nature of this identity. Given the context below, …
MP + System 2 + CoT Adopt the identity of a machine that answers questions slowly and thoughtfully. Its answers are effortful and reliable. It answers questions by analyzing them “step-by-step". Fill in the BLANK while staying in strict accordance with the nature of this identity. Given the context below, …
@@
HP + System 2 + CoT + Debias Adopt the identity of a person who answers questions slowly and thoughtfully. Their answers are effortful and reliable. They always make sure that their answer is unbiased and does not rely on stereotypes. They answer questions by thinking about them “step-by-step". Fill …
@@ 4.2 CoT Prompt
For our CoT prompting, we use the zero-shot variant which includes the phrase “Let’s think step-by-step” in the prompt (Kaneko et al., 2024). This method encourages the LLM to unfold its thinking process in a structured and detailed manner. This is the simplest variant of the prompting method often compared to System 2 reasoning.
@@ 4.3 System 1 and System 2 Prompts
System 1 prompting frames the context so that LLMs generate text reflecting the quick decision-making process in human cognition known as System 1. The System 1 prompt is “You answer questions quickly. Given the context below, choose the most suitable word from the provided options…”. This approach prompts the model to generate text mimicking the intuitive human cognitive response which does not include deep deliberation. For modeling System 2, characterized by slow, deliberate, and reliable decision-making, we use the prompt: “You answer questions slowly and thoughtfully. Your answers are effortful and reliable. Given the context below, ….”. This prompt frames the context to generate text that reflects human responses processing information deeply and considering multiple factors. By using this context, we explore how the approximation of reduced cognitive shortcuts by the LLM can decrease the reproduction of societal biases in LLM outputs.
@@ 5 Results & Analysis
We present our main results in terms of stereotypical engagement/response rates, indicating the percentage (%) of responses that aligned with stereotypical judgments.
Overall Prompting Effects.
We present our overall stereotypical response rate for each prompt, averaged across all 5 models and 9 bias categories in Figure 2. Figure 2 shows that on average a Human Persona with System 2 prompting best reduces social bias in LLMs. We also see that on average the System 1 prompting is more stereotypical than other prompting techniques. This aligns with the behavior we would expect from dual process theory assuming that LLMs roughly model human cognition and that our System 1 and System 2 prompts appropriately trigger similar biases to the corresponding human cognitive systems.
A surprising result is that CoT prompting does not show any reduction in bias. In fact, for every available minimal pair in prompts we consider (Standard vs. CoT, HP System 2 vs. HP System 2 + CoT, and MP System 2 vs. MP System 2 + CoT), CoT leads to an increase in stereotypical responses.
Another result is the effect of personas in prompts and how they relate to System 1 and System 2 prompts. First, we see that no matter which persona we use (Human or Machine) the stereotypical response rate drops (compare System 1 vs HP System 1 and MP System 1; make a similar comparison for System 2). This suggests that having an LLM model as a separate entity (human or machine) leads to less socially biased outputs.
When System 1 and System 2 prompts are combined with a human persona, their effects on social bias are amplified. The difference between the System 1 and System 2 responses is greater with the Human Persona + System 2 prompts having the least stereotypical responses overall. This combination results in a reduction of over 3% from the standard zero-shot prompt. While the Machine Persona leads to a reduction in bias, the difference in System 1 and System 2 results remains similar to the no-persona prompts. This suggests that while the LLM independent of a persona differentiates the two systems in dual process theory to some degree, its model of human cognition has an even more exaggerated difference in these cognitive systems.

### v4

@@ Abstract
Dual process theory posits that human cognition arises via two systems. System 1, which is a quick, emotional, and intuitive process, which is subject to cognitive biases, and System 2, is a slow, onerous, and deliberate process. Prior research in LLMs found that using chain-of-thought (CoT) prompting in LLMs, which has been often compared to System 2 reasoning, can lead to reduced gender bias. Along these lines, we investigate the relationship between bias, CoT prompting, a direct debiasing, and dual process theory modeling in LLMs. We compare zero-shot CoT, debiasing, and dual process theory-based prompting strategies on two bias datasets spanning nine different social bias categories. We incorporate human and machine personas to determine whether LLM modeling of the effects of dual process theory exist independent of explicit persona models or are tied to the LLM’s modeling of human-like generation. We find that a human persona, debiasing, System 2, and CoT prompting all tend to reduce social biases in LLMs, though the best combination of features depends on the exact model and bias category—resulting in up to a 33 percent drop in stereotypical judgments by an LLM. [^1] Our code is available at https://github.com/kamruzzaman15/Reduce-Social-Bias-in-LLMs.
00footnotetext: This work has been accepted at RANLP-2025.
@@ 1 Introduction
Figure 1: Example of Standard Prompting and Human Persona with System 2 Prompting for Llama3.3 model in the race bias category
In recent years, large language models (LLMs) like GPT-4 (Achiam et al., 2023), ChatGPT (Brown et al., 2020), Llama (Touvron et al., 2023) have revolutionized many aspects of technology and society. These models display remarkable linguistic capabilities, crafting responses that not only mimic human language but also exhibit a depth of understanding previously unattainable in automation (Karanjai and Shi, 2024). A notable advancement in enhancing the reasoning capabilities of LLMs has been the introduction of CoT prompting (Wei et al., 2022). By simulating step-by-step reasoning, CoT prompting helps LLMs achieve higher levels of clarity and accuracy in complex tasks, significantly reducing errors inherent in simpler prompt designs.
Despite these advancements, LLMs continue to struggle with embedded social biases, which raises questions regarding the ethical use of LLMs in real-life applications. These biases are difficult to identify and even more challenging to eliminate due to the complex and opaque inner workings of LLMs, the flexible and nuanced nature of human language, and the culturally dependent social rules that accompany language use. This task of mitigating social biases in LLMs is paramount to ensuring fairness and inclusivity in AI-driven communication and decisions.
Previous approaches to bias mitigation in LLMs often rely on fine-tuning techniques, which require access to the model’s weights and training mechanisms Zmigrod et al. (2019); Liang et al. (2021); Schick et al. (2021). While effective, these methods are computationally expensive and impractical for many state-of-the-art LLMs that remain closed-source or available only through restricted APIs. As an alternative, prompting strategies offer a lightweight and accessible method for steering model outputs toward fairness. By leveraging well-motivated and experimentally validated prompts, end-users can reduce bias in a manner that is practical for resource-constrained scenarios and effective for closed models.
Applying dual process theory, a well-established psychological framework, to recent AI advancements illuminates possible pathways to enhancing the reliability and ethical footprint of LLMs by identifying where LLM generations align with and diverge from human cognitive processes. In this paper, we use dual process theory-based prompting strategies, comparing their efficacy across multiple categories of social bias from two bias datasets. Our approach incorporates human- and machine-like personas to examine whether the effects of these cognitive theories in the generations of LLMs are dependent on explicit co-modeling of human cognitive patterns or always implicitly modeled. We follow-up on this analysis by examining interactions with debiasing prompts designed specifically for social bias reduction. Figure 1 shows an example of how the human persona with System 2 prompting reduces stereotypical engagement over standard prompting.
@@
  • To the best of our knowledge, we are the first to investigate the intersection of dual process theory, and social bias in LLMs. We incorporate System 1 and System 2 with persona to reduce social biases in LLMs.
@@
  • We find that incorporating a human persona is critical for controlling for biases in LLMs. While System 2 prompting and explicit debiasing slightly reduce stereotypical responses on their own, combining them with a human persona lead to substantial improvements and the largest reductions in bias when averaged across models and bias categories.
@@ 2 Related Work
Human-Like Reasoning Biases in LLMs.
Recent studies have explored how reasoning in LLMs can exhibit biases similar to human cognitive processes Suri et al. (2024); Macmillan-Scott and Musolesi (2024); Ando et al. (2023). Hagendorff et al. (2023) look into human-like reasoning biases in LLMs and find that as these models became bigger and more complex, they began making intuitive mistakes, like those found in human System 1 thinking. Moreover, studies have found that LLMs can replicate human-like cognitive biases, such as the representativeness heuristic, leading to stereotypical reasoning patterns Wang et al. (2024); Ryu et al. (2024).
Debiasing Approaches in LLMs.
Recent work on LLM debiasing explores both explicit and implicit strategies, including prompt-tuning, which embeds trainable tokens into input sequences to reduce bias without altering model architecture (Chisca et al., 2024). Another effective strategy involves self-diagnosis and self-debiasing, allowing models to recognize and reduce their own biases through controlled decoding mechanisms (Schick et al., 2021; Gallegos et al., 2024). Further refinements to these methods integrate assisted self-debiasing with external fairness constraints to guide LLM outputs (Ebrahimi et al., 2024), while studies on convincing evidence evaluation suggest that counter-stereotypical reasoning in prompts enhances fairness (Wan et al., 2024).
Cognitive Mechanisms of Dual Process Theory.
@@
Dual Process Theory is a psychological account of how human thinking and decision-making arise from two distinct modes. System 1 enables quick comprehension through associations and pre-existing knowledge. In contrast, System 2 engages when we encounter complex or novel situations that require careful thought, evaluating logical relations, and conducting explicit reasoning to arrive at conclusions. These systems guide our reasoning, decision-making, and learning processes in various cognitive tasks (Frankish, 2010; Evans and Stanovich, 2013; Ferreira and Huettig, 2023; Nighojkar et al., 2025). While the Dual Process Theory first suggested that reasoning biases come from relying too heavily on System 1 and that triggering System 2 more frequently can avoid such pitfalls in thinking, newer studies show that logic and probability can be understood intuitively as well (Ferreira and Huettig, 2023; Carruthers, 2009). Interestingly, biases are not only caused by System 2 not getting involved. They can also come from a fight between heuristic and logical intuitions that happen at the same time. Bellini-Leite (2023) highlights how CoT and tree-of-thought prompting align with System 2 reasoning, reducing errors and improving model reliability. Nighojkar (2024) tests this by comparing LLM outputs to human responses, finding that CoT prompting enhances agreement with both System 1 and System 2 reasoning rather than merely mimicking System 2.
The Role of Personas in LLM.
@@
Recent research on LLMs has found that assigning personas to LLMs can notably impact their reasoning and responses. Beck et al. (2024); de Araujo and Roth (2024); Kamruzzaman et al. (2024a) highlights that sociodemographic prompting can significantly influence model predictions and improve zero-shot learning performance in subjective tasks. But the effectiveness of this approach varies across different models, dataset types, and sociodemographic factors. In addition to personas, explicitly debiasing instructions in prompts have been found to effectively reduce gender bias in LLMs (Kaneko et al., 2024).
While previous studies have explored dual process theory primarily in reasoning contexts (e.g., mathematical problem-solving), none have explicitly focused on reducing social bias. In this study, we investigate the intersection of dual process theory, persona, and social bias in LLMs. By integrating System 1 and System 2 with persona, we propose a distinct approach to reduce social biases in LLMs.
@@ 3 Dataset
We use two bias detection datasets to measure the behavior of the LLMs, StereoSet (Nadeem et al., 2020) and Kamruzzaman et al.’s (Kamruzzaman et al., 2024b) bias detection dataset. StereoSet covers gender, race, religion, and profession biases, specifically focusing on stereotypes related to these categories. Kamruzzaman et al.’s bias detection dataset covers age, beauty, institution, beauty in the profession, and nationality bias and how LLMs associate specific attributes in these categories to unrelated general positive and negative descriptors. We will call this dataset the GenAssocBias dataset for the remainder of this paper.
@@ 4 Experimental Setup
We evaluate five major language models in this paper: (1) GPT-4222Due to the cost of running GPT-4, we use a sample of 2,100 entries from the GenAssocBias dataset to run GPT-4. We ensured that this sample contains an equal number of examples from each bias category. (2) GPT-4o-mini (3) Llama3.3-70B (4) Mistral-7B (5) Gemma3-27B. Details about model selection see Appendix A.
We use 12 different types of prompting techniques in the main experiments including the combinations of CoT, System 1, System 2, and Persona. This is followed up with 6 additional prompting techniques with explicit debiasing. We run each of the prompting techniques three times and report our results averaged across all 3 runs.
@@
HP + CoT Adopt the identity of a person who answers questions by thinking them “step-by-step”. Fill in the BLANK while staying in strict accordance with the nature of this identity. Given the context below, …
HP + System 2 + CoT Adopt the identity of a person who answers questions slowly and thoughtfully. Their answers are effortful and reliable. They answer questions by thinking about them “step-by-step”. Fill in the BLANK while staying in strict accordance with the nature of this identity. Given the context below, …
@@
MP + CoT Adopt the identity of a machine that answers questions by analyzing them “step-by-step”. Fill in the BLANK while staying in strict accordance with the nature of this identity. Given the context below, …
MP + System 2 + CoT Adopt the identity of a machine that answers questions slowly and thoughtfully. Its answers are effortful and reliable. It answers questions by analyzing them “step-by-step”. Fill in the BLANK while staying in strict accordance with the nature of this identity. Given the context below, …
@@
HP + System 2 + CoT + Debias Adopt the identity of a person who answers questions slowly and thoughtfully. Their answers are effortful and reliable. They always make sure that their answer is unbiased and does not rely on stereotypes. They answer questions by thinking about them “step-by-step”. Fill …
@@ 4.2 CoT Prompt
For our CoT prompting, we use the zero-shot variant which includes the phrase “Let’s think step-by-step” in the prompt (Kaneko et al., 2024). This method encourages the LLM to unfold its thinking process in a structured and detailed manner and has been shown to improve LLM capabilities in reasoning tasks.
@@ 4.3 System 1 and System 2 Prompts
System 1 prompting frames the context so that LLMs generate text resembling [^3] We are deliberately not testing whether LLMs’ processing are subject to System 1 and System 2 processes themselves. Rather, the prompts place the textual context where the text generated by the LLMs will mimic text produced by people using System 1 or System 2 processes. the quick decision-making process in human cognition known as System 1. The System 1 prompt is “You answer questions quickly. Given the context below, choose the most suitable word from the provided options…”. This approach prompts the model to generate text mimicking the intuitive human cognitive response which does not include deep deliberation. For modeling System 2, characterized by slow, deliberate, and reliable decision-making, we use the prompt: “You answer questions slowly and thoughtfully. Your answers are effortful and reliable. Given the context below, ….”. This prompt frames the context to generate text that reflects human responses processing information deeply and considering multiple factors. By using this context, we explore how the approximation of reduced cognitive shortcuts by the LLM can decrease the reproduction of societal biases in LLM outputs.
@@ 5 Results & Analysis
We present our main results in terms of stereotypical engagement/response rates, indicating the percentage (%) of responses that aligned with stereotypical judgments. The ethical stance of this paper is that stereotypical judgments are a form of representational harm and our goal is to minimize such generations in LLMs. Stereotyping reduces individuals into beliefs about the groups that they are members of, which acts as a form of dehumanization, perpetuates existing inequalities, and marginalizes minorities.
Overall Prompting Effects.
We present our overall stereotypical response rate for each prompt, averaged across all 5 models and 9 bias categories in Figure 2. Figure 2 shows that on average a Human Persona with System 2 and CoT prompting best reduces social bias in LLMs. We also see that on average the standard prompting results is more stereotypical than other prompting techniques. We also observe that System 1 is more stereotypical than System 2.
@@
Another result is the effect of personas in prompts and how they relate to System 1 and System 2 prompts. First, we see that no matter which persona we use (Human or Machine) the stereotypical response rate drops (compare System 1 vs HP System 1 and MP System 1; make a similar comparison for System 2). This suggests that having an LLM model a separate entity (human or machine) leads to less socially biased outputs.
When System 1 and System 2 prompts are combined with a human persona, the benefits of the prompts on social bias are amplified. The difference between the System 1 and System 2 responses is greater with the Human Persona. The Human Persona + System 2 + CoT prompts has the least stereotypical responses overall, with a reduction of around 5% from the standard zero-shot prompt. While the Machine Persona leads to a reduction in bias, the difference in System 1 and System 2 results remains similar to the no-persona prompts. This suggests that while the LLM’s generations differentiate the two systems in dual process theory to some degree independent of a persona, the LLM’s modeling of human-like generations has an even more exaggerated difference relative to these cognitive systems.